Q24): Provide one example of a complete molecule with a functional group that has a carbon at the following oxidation states: -3,-2, -1, 0, +1, +2,+ 3, +4. Please do not use R groups as symbols for substituents, but instead use actual substituents.

Q25) Provide TWO additional molecular examples for functional groups that have carbon atoms at the oxidation +2 and +3 states, beyond the one provide in Q24.

needs to be some type of carbohydrate, (poly)sacharide, or sugar. thanks
Answer #1

#24 and #25:

In organic compounds, Oxidation state of H attached to more electronegative elements like C, O, N etc is -1.

Oxidation state of O, S is -2.

Oxidation state of halogens like F, Cl, Br and I is -1.

Oxidation state of N is -3.

When two same C atoms are attached, the oxidation state of each C due to that bond is 0.

#(a): -3 oxidation state of C:

To get -3 oxidation state of C, it should be attached to 3 H(=3*1 = 3) and a C (=0) atom.

Hence the molecule is ethane, CH3 - CH3.

In CH3-CH3, the oxidation state of each C atom is -3.

#(b): -2 oxidation state of C:

To -2 oxidation state, the C atom should be attached to 2H atoms and C atom. This is possible in case of alkene. Hence in ethene, CH2=CH2, the oxidation state of each C atom is -2.

-2 oxidation state is also found in the C of CH3-OH. In CH3-OH, the C has oxidation state of -3 due to 3H and +1 due to O atom. Hence total oxidation state of C in CH3OH is -3+1 = -2.

Hence the answers are: ethene, CH2=CH2 and CH3-OH

#(c): -1 oxidation state of C:

To -1 oxidation state, the C atom should be attached to 1H atom and C atom. This is possible in case of alkyne. Hence in ethyne, CHCH, the oxidation state of each C atom is -1.

Hence the answer is ethyne, CHCH.

#(d): 0 oxidation state of C:

In case of 0 oxidation state, the C atom should contains equal amount of + and - partial oxidation state. In methanal, HCHO, the oxidation state of C due 2H is -2 and due to O(=O) is +2. Hence total oxidation state of C = -2+2 = 0.

Hence the answer is methanal, HCHO.

#(e): +1 oxidation state of C:

In CH3-CHO, the oxidation state of C in -CHO is +1.

Hence the answer is ethanal, CH3CHO.

#(f): +2 oxidation state of C:

The oxidation state of C in methanoic acid, HCOOH is +2

Also the oxidation state of C=O carbon in acetone, CH3COCH3 is +2

Hence the answers are: methanoic acid, HCOOH and acetone, CH3COCH3 (Only the C=O C has +2 oxidation state).

#(g): +3 oxidation state of C:

The oxidation state of COOH carbon in ethanoic acid, CH3COOH, propanoic acid, CH3CH2COOH etc. have +3 oxidation state.

Hence the answers are: ethanoic acid, CH3COOH, propanoic acid, CH3CH2COOH (only the COOH carbon).

#(h): +4 oxidation state of C:

Oxidation state of C in CCl4 is +4. C has +1 partial oxidation state due to each Cl making a total of +4.

Hence the answer is CCl4